### v2

# MedCare: Advancing Medical LLMs through Decoupling Clinical Alignment and Knowledge Aggregation

Abstract
Large language models (LLMs) have shown substantial progress in natural language understanding and generation, proving valuable especially in the medical field. Despite advancements, challenges persist due to the complexity and diversity inherent in medical tasks, which can be categorized as knowledge-intensive tasks and alignment-required tasks. Previous approaches either ignore the latter task or focus on a minority of tasks and hence lose generalization. To address these drawbacks, we propose a progressive fine-tuning pipeline. This pipeline employs a Knowledge Aggregator and a Noise Aggregator to encode diverse knowledge in the first stage and filter out detrimental information. In the second stage, we drop the Noise Aggregator to avoid the interference of suboptimal representation and leverage an additional alignment module optimized towards an orthogonal direction to the knowledge space to mitigate knowledge forgetting. Based on this two-stage paradigm, we proposed a Medical LLM through decoupling Clinical Alignment and Knowledge Aggregation (MedCare), which is designed to achieve state-of-the-art (SOTA) performance on over 20 medical tasks, as well as SOTA results on specific medical alignment tasks. Various model sizes of MedCare (1.8B, 7B, 14B) all demonstrate significant improvements over existing models with similar model sizes.
MedCare: Advancing Medical LLMs through Decoupling Clinical Alignment and Knowledge Aggregation
Yusheng Liaothanks: Equal contribution.,1, Shuyang Jiang∗,2,3, Yanfeng Wang1,3, Yu Wangthanks: Corresponding Author,1,3 1Shanghai Jiao Tong University 2Fudan University 2Shanghai Artificial Intelligence Laboratory {liao20160907,,wangyanfeng622,yuwangsjtu}@sjtu.edu.cn shuyangjiang23@m.fudan.edu.cn
1 Introduction
Large Language Models (LLMs) (OpenAI, 2022; AI, 2024; Singhal et al., 2023a) have made significant strides in the realms of natural language generation and thereby finding extensive applications across a broad spectrum of disciplines (Liu et al., 2023a; Deng et al., 2024). Among these, the medical field has attracted considerable attention due to its importance and demand. Much effort has been dedicated to researching and developing specialized LLMs for this domain (Li et al., 2023b; Wang et al., 2023a; Singhal et al., 2023a).
Figure 1: Examples of two types of medical tasks. Knowledge-intensive tasks require models to possess sufficient knowledge, whereas alignment-required tasks additionally necessitate the model to meet specific requirements criteria.
Although progress has been made, LLMs still face challenges due to the complexity and diversity of medical tasks. This is particularly evident when data is intentionally structured to represent a diverse set of tasks (Chung et al., 2022; Wei et al., 2021; Sanh et al., 2021). We categorize medical tasks into two types: knowledge-intensive tasks and alignment-required tasks. Knowledge-intensive tasks, such as medical question answering (Zhang et al., 2018) and medical dialogues (Liu et al., 2022a; Zhao et al., 2022), primarily test the model’s understanding of internal medical knowledge. Conversely, alignment-required tasks, like clinical terminology standardization (Zhang et al., 2022b) and medical entity recognition (Hongying et al., 2021), demand not only medical knowledge but also strict adherence to output formats. Figure 1 illustrates these two categories.
However, previous medical LLMs, such as HuatuoGPT (Wang et al., 2023a), HuatuoGPT-II (Chen et al., 2023a), Med-PaLM (Singhal et al., 2023a) and BioMistral (Labrak et al., 2024), have primarily focused on enhancing the encoding of medical knowledge in LLMs, neglecting the requirements of alignment-required tasks Van Veen et al. (2024). On the other hand, alignment-oriented methods Liu et al. (2023b); Cai et al. (2023); Wang et al. (2023a) that perform alignment through fine-tuning often suffer from an "alignment tax". This results in knowledge forgetting and performance drops in knowledge-intensive tasks (Lin et al., 2023; Gekhman et al., 2024). These issues limit the practical use of LLMs in healthcare.
To create a more practical medical LLM, in this paper we propose a novel training pipeline consisting of two progressive fine-tuning stages: miscellaneous knowledge aggregation (MKA) and downstream alignment (DA). In the first stage, we introduce two modules, Knowledge Aggregator (KA) and Noise Aggregator (NA), to encode advantageous knowledge and noisy contents, respectively. The Noise Aggregator is asymmetric with respect to Knowledge Aggregator, which explicitly induces Knowledge Aggregator to perform multi-task knowledge extraction. To avoid catastrophically parameterized knowledge disruption, we innovatively remove the updated Noise Aggregator but retain only the Knowledge Aggregator after training. Following this stage, we introduce an alignment module to cater to the downstream alignment requirements from the specific task. We add an orthogonal regularization term to ensure non-overlapping between the optimization space of alignment and knowledge space in the first stage. Built upon this pipeline, we propose MedCare with three sizes (1.8B, 7B, 14B), a specialized LLM tailored for both knowledge-intensive tasks and alignment-required tasks, derived from the Qwen1.5 series. Figure 2 provides a visual representation of MedCare.
In summary, our contributions are as follows:
  • We introduce a taxonomy for medical tasks, which divides all tasks into knowledge-intensive tasks and alignment-required tasks. This taxonomy reinforces the practical usability requirements for medical LLMs.
  • We introduce a two-stage fine-tuning pipeline that balances knowledge maintenance and downstream alignment. This approach not only encodes knowledge without disruption but also adapts to specific tasks with minimal knowledge forgetting.
  • Based on the progressive pipeline, we introduce MedCare, a medical LLM to simultaneously encode massive knowledge and align with practical requirements under the medical multi-task taxonomy. To our knowledge, MedCare is the first LLM to effectively handle such a wide spectrum of tasks in the medical domain with few alignment taxes.
  • We conduct extensive experiments on over 20 knowledge-intensive and alignment-required tasks, benchmarking MedCare against a diverse array of established language models. Our results demonstrate the superior performance of MedCare, affirming its effectiveness for both genres of tasks.
Figure 2: Overview of the proposed MedCare. In the MKA stage, MedCare encodes advantageous knowledge and noisy contents with Knowledge Aggregator and Noise Aggregator from both types of tasks, respectively. The updated Noise Aggregator is removed to avoid the knowledge disruption. In the DA stage, an additional alignment module and orthogonal regularization are introduced to cater to the requirements of the alignment tasks.
2 Preliminaries
Low-rank Adaptation (LoRA)
Given an input sequence $\mathbf{X}\in\mathbb{R}^{m\times d}$, LoRA (Hu et al., 2021) proves that the update of original linear layers $\Delta W$ in large language models is of low-rank, and can be decomposed into the multiplication of two compact matrices $AB$. The pretrained weight $W$ is frozen in the training phase and does not undergo gradient updates, while $A$ and $B$ are trainable parameters and contribute together to the forward pass:
$\mathbf{H}=W\mathbf{X}+\mathbf{X}\Delta W=\mathbf{X}W+\frac{\alpha}{r}\mathbf{X}AB$ (1)
where $W\in\mathbb{R}^{d\times d^{\prime}},A\in\mathbb{R}^{d\times r},B\in\mathbb{R}^{r\times d^{\prime}}$ and $r\ll d,d^{\prime}$. $\alpha$ is the scaling factor. Without loss of generality, we omit the layer index for the following formula. At the beginning of training, $B$ is initialized to an all-zero matrix and $A$ uses Gaussian initialization to make sure that the product $AB$ is zero at initialization.
Mixture of LoRA
The mixture of the LoRA (MoLoRA) module is based on the Feed-forward Network (FFN) module in the LLMs since experts of FFN can store diverse task knowledge (Geva et al., 2020) from multi-task learning. In the popular LLaMA-style FFN, the input $\mathbf{X}$ is computed using an SiLU (Elfwing et al., 2018) gate as follow:
$\mathbf{X}=(\mathbf{X}W_{u}\cdot\mathrm{silu}(\mathbf{X}W_{g}))W_{d}$ (2)
where $W_{g}\in\mathbb{R}^{d^{\prime}\times d},W_{u}\in\mathbb{R}^{d^{\prime}\times d},W_{d}\in\mathbb{R}^{d\times d^{\prime}}$ and $d^{\prime}=8/3d$. In the LoRA forward passes, each linear layer $W\in\{W_{d},W_{u},W_{g}\}$ is updated using the LoRA module described in Eq.1. Consider an MoLoRA module with $E$ experts, and each expert is denoted as $\{E_{i}\}_{i=1}^{E}$, the forward pass of the linear layer is formulated as:
$\mathbf{H}$ $=\mathbf{X}W+\mathbf{X}\Delta W$ (3)
$\mathbf{X}\Delta W$ $=\sum_{i=1}^{K}G(\mathbf{X})_{i}E_{i}(\mathbf{X})$ (4)
where $\mathbf{H}$ is the processed output and $G(\cdot)=\mathrm{softmax}(\mathbf{X}W_{r}),W_{r}\in\mathbb{R}^{d\times E}$ is the router in MoLoRA with top-$K$ selection which holds the top-$K$ affinity with respect to current input $\mathbf{X}$. Assuming each LoRA expert shares the same rank $r$ and $\alpha$ and we obtain the weight $w_{i}$ from the router $G(\cdot)$ as $w_{i}=G(\mathbf{X})_{i}$, the overall output of the MoLoRA linear layer can be formulated as:
$\mathbf{h}=W\mathbf{x}+\frac{\alpha}{r}\sum_{i=1}^{K}w_{i}\cdot A_{i}B_{i}\mathbf{x}$ (5)
where $\mathbf{x},\mathbf{h}\in\mathbb{R}^{d}$ is any token representation of $\mathbf{X}$ and the frozen $W$ inherits from the pre-trained linear weight.
3 MedCare
In this paper, we categorize the medical task into two primary genres: knowledge-intensive tasks and alignment-required tasks. Knowledge-intensive tasks require LLMs to seamlessly transfer both medical knowledge and their general domain reasoning capabilities into the medical domain. Models leverage the internal broad pre-trained knowledge without significantly disrupting the pretrained distribution, to fulfill domain-specific medical reasoning tasks. In contrast, alignment-required tasks pose a greater challenge. These tasks typically deviate more substantially from the model’s original training paradigms, which can lead to hallucinations (Gekhman et al., 2024) and the aliasing of pre-trained knowledge (Dou et al., 2023). To address this challenge, we propose MedCare, a method designed to simultaneously acquire the necessary knowledge for the general adaptation of LLMs and to achieve specialized alignment for specific tasks. The fine-tuning process of MedCare comprises two stages: miscellaneous knowledge aggregation (MKA; §3.2) and downstream alignment (DA; §3.3). The training procedure for MedCare is illustrated in Figure 2.
3.1 Problem Formulation
In the MKA stage, the dataset $\mathcal{D}_{\text{MKA}}$ comprises $N$ samples $\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$, derived from tasks that are both knowledge-intensive and require alignment. Here, $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$ represent the input and target for the models, respectively. LLMs assimilate diverse knowledge from these tasks to augment their knowledge and reasoning capabilities. Subsequently, these LLMs are exposed to a specific subset of the alignment-required data $\mathcal{D}_{\rm DA}$, which is necessary for learning the intricacies of various alignment-required tasks, including named entity recognition with specificity and report analysis that demands explanatory detail.
3.2 Miscellaneous Knowledge Aggregation
In this stage, the models acquire common medical knowledge from both types of tasks. To absorb miscellaneous knowledge into single LLMs without interference, we adopt the MoLoRA structure (Li et al., 2024; Su et al., 2024; Feng et al., 2024; Luo et al., 2024) as the Noise Aggregator and introduce shared experts as the Knowledge Aggregator to further circumvent low generalization due to the high specialization of each expert (Gou et al., 2023) and parameter redundancy (Dai et al., 2024) on the FFN module of the models. Through backpropagation, the KA acquires common knowledge from multiple datasets, while NA learns the distinct alignment-required tasks. Consequently, the KA absorbs common knowledge encoded in each task, while the NA learns the interfering alignment requirements.
$\mathbf{X}$ $=\mathbf{X}W+\mathrm{KA}(\mathbf{X})+\mathrm{NA}(\mathbf{X})$ (6)
$=\mathbf{X}W+\mathrm{LoRA}_{s}(\mathbf{X})+\mathrm{MoLoRA}(\mathbf{X})$ (7)
where $W$ denotes any weight matrix in FFN modules. The rank of NA is $r$, and the rank of KA is $r^{\prime}=sr$, where $s$ is the number of the share experts. For the attention module, we only add a vanilla LoRA module to each projection layer $W$ as Eq.1 to guarantee sufficient optimization space. As most task knowledge is encoded in the common knowledge aggregator, we discard the separate knowledge aggregator after this training stage. In other words, we introduce redundant parameters to avoid erroneous leverage of task-agnostic knowledge through decoupling the KA and NA modules, employing only the KA module:
$\mathbf{X}=\mathbf{X}W+\mathrm{KA}(\mathbf{X})$ (8)
While discarding the separate KA prevents the model’s knowledge from being disturbed, it also compromises the model’s alignment capabilities, thus leading to suboptimal performance on alignment-required tasks. Therefore, a second stage of training is necessary to enable the model to learn the requirements of alignment tasks while preserving its knowledge reasoning capabilities.
Model MedQA MMB. CMB CMExam CMMLU CEval PLE Pha PLE TCM Avg.
Llama3-8B 59.40 63.78 41.63 44.99 51.40 53.66 38.33 33.54 48.34
ChatGLM3-6B 44.51 51.34 39.81 43.21 46.97 48.80 34.60 32.90 42.77
Baichuan2-7B 45.97 52.39 46.33 50.48 50.74 51.47 44.60 42.10 48.01
Baichuan2-13B 49.42 56.71 50.87 54.90 52.95 58.67 44.20 41.70 51.18
Qwen1.5-7B 74.46 78.58 61.33 58.24 66.77 68.29 32.29 29.17 58.64
Qwen1.5-14B 81.93 84.33 64.33 57.79 68.76 78.05 48.33 38.33 65.23
ChatGPT 37.51 40.08 43.26 46.51 50.37 48.80 41.20 31.20 42.81
HuatuoGPT-II-7B 59.22 62.03 60.39 65.81 59.08 62.40 47.70 47.50 58.02
HuatuoGPT-II-13B 75.77 78.69 63.34 68.98 61.45 64.00 52.90 51.60 64.59
MedCare-1.8B 56.80 65.62 43.04 47.83 48.82 51.22 38.13 35.21 48.33
      $\mapsto$ w/o DA 61.70 64.62 49.28 52.21 51.18 60.98 36.04 39.58 51.95
MedCare-7B 76.77 80.79 60.13 65.33 64.33 65.85 47.29 52.08 64.07
      $\mapsto$ w/o DA 75.16 79.36 61.60 66.85 66.25 70.73 50.83 53.33 65.51
MedCare-14B 81.44 84.76 64.17 68.74 71.57 78.05 54.17 54.58 69.69
      $\mapsto$ w/o DA 77.41 83.36 66.79 71.77 69.42 82.93 57.29 54.17 70.39
Table 1: The results on medical knowledge exams. The results of the MedQA are evaluated on the Chinese Mainland subset. ‘MMB.’ indicates the Chinese subset of the MMedBench. ‘PLE Pha’ and ‘PLE TCM’ indicate the Pharmacy and Traditional Chinese Medicine tracks of the 2023 Chinese National Pharmacist Licensure Examination. Note that for the general benchmarks, CMMLU and CEval, we only chose questions related to the medical domain.
3.3 Downstream Alignment
In this stage, MedCare merges back the LoRA weights of self-attention into the pretrained model and freezes the self-attention module to ensure the instruction-following ability of LLMs (Wu et al., 2023). We introduce an additional alignment LoRA module $\mathrm{Align}(\cdot)$ in the FFN module to acquire specific alignment knowledge from the alignment dataset, the forward pass of linear layers of FFN can be formulated as:
$\mathbf{X}$ $=\mathbf{X}W+\mathrm{KA}(\mathbf{X})+\mathrm{Align}(\mathbf{X})$ (9)
Following Wang et al. (2023c), we introduce an orthogonal loss to ensure that the new alignment task is learned in a direction orthogonal to the original knowledge task. For the LoRA weights of KA $\{A^{p}_{k},B^{p}_{k}\}_{p=1}^{P}$ and Align $\{A^{p}_{d},B^{p}_{d}\}_{p=1}^{P}$, where $P$ is the total number of modules, we achieve the orthogonal subspaces of the alignment stage as:
$\arg\min O^{p}_{k,d}={A^{p}_{k}}^{\top}A^{p}_{d}$ (10)
Based on the subspace learning, the learning objective of this stage is expressed as
$\mathcal{L}$ $=\mathcal{L}_{\mathrm{nll}}+\lambda\mathcal{L}_{\mathrm{orth}}$ (11)
$\abs\displaystyle=-\sum_{\mathbf{x},\mathbf{y}\in\mathcal{D}_{\mathrm{DA}}}\log p_{\theta}(\mathbf{y}|\mathbf{x})+\lambda\sum_{p=1}^{P}\abs$
[…]
$W^{\prime}=W+\frac{\alpha}{r^{\prime}}A_{k}B_{k}+\frac{\alpha}{r^{\prime}}A_{d}B_{d}$ (12)
4 Experiments
MedCare is built upon the Qwen1.5-Chat  [^1] https://github.com/QwenLM/Qwen1.5 series. For knowledge-intensive tasks, we adopt the Chinese Mainland test set of MedQA (Jin et al., 2021), the Chinese subset of MMedBench (Qiu et al., 2024), and two comprehensive Chinese medical exam datasets, CMB (Wang et al., 2023d) and CMExam (Liu et al., 2024a). We also collect the medical parts of the general benchmarks, including CMMLU (Li et al., 2023a) and CEval (Huang et al., 2023). Besides, we test the performance of the models on the flash exam questions from the 2023 Chinese National Pharmacist Licensure Examination (PLE) collected by Chen et al. (2023a). For the alignment-required tasks, we use CBLUE (Zhang et al., 2022a) and our newly developed Chinese Clinical Task Evaluation (CCTE) dataset. More details of the experiment settings can be found in Appendix B.
Figure 3: Results on 16 tasks in CBLUE. ChatGLM-6B is fine-tuned on the CBLUE dataset with LoRA and ChatGPT is augmented by in-context learning. The results are obtained from the official implementation of CBLUE.
Report Generation Image Analysis Discharge Instruction Examination Education
Model Flu. Rel. Com. Pro. Avg. Flu. Rel. Com. Pro. Avg. Flu. Rel. Com. Pro. Avg. Flu. Rel. Com. Pro. Avg. Avg.
ChatGLM2-6B 4.16 2.66 2.64 2.52 3.00 4.88 4.06 4.04 3.70 4.17 4.98 3.46 3.12 3.24 3.70 3.68 2.94 2.44 2.76 2.96 3.46
ChatGLM3-6B 4.52 2.78 2.66 2.52 3.12 4.98 4.16 4.20 3.94 4.32 5.00 3.34 3.24 3.04 3.66 4.52 2.88 2.14 2.56 3.03 3.53
Baichuan2-7B 4.88 3.36 3.30 3.40 3.74 5.00 4.40 4.40 4.02 4.46 5.00 3.96 3.84 3.82 4.16 3.98 2.74 1.68 2.20 2.65 3.75
Baichuan2-13B 4.94 3.42 3.40 3.46 3.81 5.00 4.42 4.42 4.22 4.52 5.00 3.92 3.92 3.88 4.18 4.56 2.88 1.40 2.18 2.76 3.81
Qwen1.5-7B 4.72 3.22 3.26 3.12 3.58 5.00 4.64 4.64 4.42 4.68 5.00 4.18 4.40 4.22 4.45 4.94 3.34 1.80 2.94 3.26 3.99
Qwen1.5-14B 4.86 3.54 3.50 3.50 3.85 5.00 4.66 4.66 4.44 4.69 5.00 4.10 4.16 4.10 4.34 4.80 3.28 1.68 2.56 3.08 3.99
ChatGPT 4.88 3.32 2.96 3.12 3.57 4.98 4.30 4.04 4.00 4.33 5.00 3.90 3.68 3.64 4.06 4.60 3.36 2.44 2.96 3.34 3.82
HuatuoGPT-II-7B 4.80 3.16 3.12 3.30 3.60 5.00 4.22 4.26 4.04 4.38 5.00 3.96 3.90 3.88 4.19 5.00 4.08 3.88 3.90 4.22 4.09
HuatuoGPT-II-13B 4.78 3.22 3.16 3.40 3.64 5.00 4.34 4.30 4.14 4.45 5.00 3.90 3.88 3.84 4.16 4.90 4.30 4.02 3.96 4.30 4.13
MedCare-1.8B 5.00 3.66 3.54 3.78 4.00 4.98 3.92 4.00 3.86 4.19 5.00 4.18 4.24 4.12 4.39 4.72 4.20 3.90 3.88 4.18 4.19
MedCare-7B 5.00 3.78 3.68 3.80 4.07 5.00 4.12 4.24 4.12 4.37 5.00 4.36 4.42 4.38 4.54 4.96 4.42 4.40 4.34 4.53 4.38
MedCare-14B 4.98 3.74 3.62 3.70 4.01 5.00 4.14 4.24 4.10 4.37 5.00 4.46 4.46 4.40 4.58 4.94 4.54 4.48 4.44 4.60 4.39
Table 2: Results on medical alignment task CCTE. ‘Flu.’ indicates ‘Fluency’, ‘Rel.’ indicates ‘Relevance’, ‘Com.’ indicates ‘Completeness’, and ‘Pro.’ indicates ‘Proficiency’. The maximum value of all scores is 5.
4.1 Results on Knowledge-Intensive Tasks
We utilized eight distinct datasets to rigorously assess the medical knowledge capabilities of the proposed model MedCare for comprehensive evaluation. As detailed in Table 1, MedCare has demonstrated exceptional proficiency in medical knowledge, particularly notable in the MedCare-14B, which has outperformed both ChatGPT and the leading open-source medical model, HuatuoGPT-II (13B). Remarkably, the MedCare-1.8B model matches the performance of ChatGLM3-6b, while MedCare-7B achieves comparable results to the HuatuoGPT-II (13B). Besides, without the Downstream Alignment, MedCare achieves better performance.
To mitigate potential biases caused by data leakage, we further evaluated our models on the latest 2023 Chinese National Pharmacist Licensure Examinations (PLE) (Chen et al., 2023a), with the detailed results shown in Table 6. This assessment confirms that MedCare continues to outpace the competition, with MedCare-14B w/o DA surpassing the performance of GPT-4. These findings underscore MedCare’s superior utilization of medical knowledge, achieving higher expertise levels with fewer parameters.
4.2 Results on Alignment-Required Tasks
To evaluate the medical alignment ability of the models, we tested CBLUE and CCTE with 20 distinct alignment-required tasks. CBLUE requires models to generate outputs in a prescribed format. A higher score indicates the model aligns more closely with the task requirements. As depicted in Figure 3, the results indicate that even when augmented by In-Context Learning (ICL), ChatGPT still fails to fulfill such specific format requirements. Besides, MedCare-7B surpasses the established CBLUE baseline ChatGLM on nearly all the tasks with an average of more than 3 points. Since CBLUE strictly requires the output format, we mainly evaluate model alignment capabilities through ablation studies discussed in §5 instead of comparing it with other LLMs.
CCTE encompasses four clinical tasks that, while not specifying the format of model outputs, require a higher level of professional expertise. For this assessment, we employed GPT-4 to score the outputs across four dimensions comprehensively. As shown in Table 2, it is evident that even the MedCare-1.8B variant outperforms both general and medical-specific large language models, further confirming MedCare ’s robust potential in clinical settings. More details about CCTE are discussed in Appendix C.
MKA DA Knowledge-Intensive Alignment-Required
KA  NA w/o $\mathcal{L}_{\mathrm{orth}}$ w/ $\mathcal{L}_{\mathrm{orth}}$ CMMLU CEval PLE Pha PLE TCM Avg. CBLUE CCTE
Base Model 48.45 53.66 31.46 25.63 39.80 3.76 3.36
46.23 41.46 33.54 33.96 38.80 51.41 4.04
50.07 45.00 36.04 33.54 41.16 57.09 4.15
[…]
45.79 41.46 37.50 34.38 39.78 56.26 4.02
[…]
4.3 Ablation Experiments
The results of the ablation experiments are shown in Table 3. For the methods only fine-tuned with the MKA stage, inference with NA demonstrated minimal improvement in knowledge examination performance. Without the disturbance of the NA, MedCare achieved the best knowledge performance but failed to complete the alignment-required tasks. For the two-stage fine-tuning methods (MKA+DA), further fine-tuned with the NA module or without orthogonal regularization led to a noticeable reduction in performance on knowledge-intensive tasks. The orthogonal regularization can avoid parameter redundancy, facilitating more effective alignment learning while mitigating the loss of reasoning capability.
Model Knowledge-Intensive Alignment-Required
CMMLU CEval PLE Pha. PLE TCM. Avg. CBLUE CCTE
Base Models 48.45 53.66 31.46 25.63 39.80 3.76 3.36
Parallel LoRA 48.67 43.90 35.00 34.17 40.43 55.05 4.09
      $\mapsto$ w/o LoRA 1 51.26 51.22 34.79 35.00 43.07 11.43 3.89
      $\mapsto$ w/o LoRA 2 50.52 48.78 35.63 33.75 42.17 11.87 3.88
MedCare 50.07 45.00 36.04 33.54 41.16 57.09 4.15
      $\mapsto$ w/o NA 51.18 60.98 36.04 39.58 46.95 8.15 3.80
      $\mapsto$ w/o KA 49.19 39.02 34.17 32.50 38.72 50.22 4.06
Table 4: Performance of partial experts with MKA fine-tuning stage. ‘Parallel LoRA’ has two identical LoRAs on FFN modules, which are numbered with LoRA1 and LoRA2 for distinguishment.
5 Discussion
In this section, we discuss the following research questions (RQ) of the proposed MedCare:
  • RQ1: Why does Noise Aggregator have a negative impact on the task performance?
  • RQ2: Are the roles of Knowledge Aggregator and Noise Aggregator determined by the model architecture?
  • RQ3: How do both two fine-tuning stages improve the model’s capabilities?
  • RQ4: Can MedCare still learn the knowledge effectively when the scale of fine-tuning corpus becomes smaller?
Figure 4: Mismatch between expert activation times and performance on CEval. (a) Activation times of each expert combination. (b) Performance of only used each expert combination. The $i$-th raw and $j$-th column indicates the combinations of the $i$-th and $j$-th experts. The accuracy of the vanilla MoLoRA inference is 45.00.
Figure 5: Performance with different sizes of alignment-required datasets for DA fine-tuning stage. (a) The average performance on knowledge-intensive tasks. (b) The average performance on CBLUE. (c) The average performance on CCTE. ‘MedCare w/o MKA’ indicates the model is fine-tuned with only the second stage. Note that the score of the knowledge examinations is the average of the CMMLU, CEval, PLE Pharamy, and PLE TCM.
Response to RQ1: Noise Aggregator interferes with models because of the routing mismatch.
We investigate the effect of different combinations of MoLoRA experts in Noise Aggregator on the final performance to demonstrate the routing mismatch. We first compute the activation frequencies for each combination of the experts in vanilla MoLoRA inference. Meanwhile, different expert combinations are designated to test the performance of the models. The results in Figure 4 show the mismatch between the experts’ activation times and their corresponding performance, which indicates that the router in the MoLoRA module failed to select the experts with optimal performance for each task and even caused the performance drop during inference. By discarding the Noise Aggregator, LLMs can learn knowledge more effectively and surpass the vanilla MoLoRA models with an average of more than 5 points, as shown in Table 4.
Response to RQ2: Yes. Symmetric structure failed to decouple the learning process into the common knowledge and task-specific requirements.
To demonstrate this, we tested the models with two parallel LoRA modules to investigate the role of the Noise Aggregator. As shown in Table 4, the parallel LoRA module failed to decouple the learning process due to its symmetric structure. Conversely, in the structure of MedCare, the Knowledge Aggregator predominantly acquires common knowledge from the fine-tuning corpus, while the Noise Aggregator focuses more on learning alignment formats.
Response to RQ3: The KA stage improves the knowledge capacity of the models and the DA stage adapts the models to learn the target formats.
We explore the respective role of each fine-tuning stage and empirically validate each merit. Figure 5 shows the results on knowledge examination, CBLUE, and CCTE, respectively. For the first stage, it is obvious that the models with the MKA fine-tuning stage not only achieve better performance on the knowledge examination tasks but also align with the format of the downstream dataset with greater ease. Although discarding Noise Aggregator after the first stage of fine-tuning reduces the alignment performance of the model, the target format can be learned faster by retraining the model with the second-stage medical adaptation. As more second-stage aligned data is added, the model’s alignment capability significantly improves without compromising its knowledge capacity. This suggests that the second stage of the fine-tuning process primarily helps the model align with the format, rather than learning new knowledge.
Response to RQ4: Yes. Even with limited data, MedCare can enhance its medical knowledge proficiency.
To show the effective knowledge learning ability of MedCare, we compare the performance of models with varying scales of the first-stage fine-tuning corpus, and present the results in Figure 6. The vanilla MoLoRA suffers from performance degradation when the size of the training-corpus is smaller than 50%, while MedCare utilizes the knowledge in the corpus among all sizes of the data and surpasses the performance of the base model consistently, demonstrating that MedCare has better knowledge utilization ability.
Figure 6: Average performance on the medical knowledge examination with different knowledge aggregation learning data size. ‘Base Model’ indicates the performance of the Qwen1.5-1.8B without fine-tuning.
6 Conclusions
In this paper, we first categorize medical tasks into knowledge-intensive tasks and alignment-required tasks to reinforce the practical usability requirements for medical LLMs. Then, we propose a two-stage fine-tuning framework to adapt LLMs to medical domains and mitigate knowledge performance degradation and propose MedCare. The experiment results show that MedCare can achieve promising performance on over 20 knowledge-intensive and alignment-required tasks.
Limitations
In this paper, we propose a two-stage fine-tuning framework that mitigates the damage and loss of pre-trained knowledge in large language models during downstream fine-tuning for medical tasks. Despite its effectiveness, our approach still exhibits limitations. Through knowledge-aggregative learning, our method enables large language models to more effectively assimilate the knowledge of the fine-tuning corpus, thereby enhancing the knowledge capabilities of the final model beyond the baseline. However, the alignment fine-tuning in the second stage still adversely affects the model’s knowledge capabilities. Furthermore, our method does not yet allow for the decoupling of knowledge and format learning directly, but instead requires two-stage training. Addressing these two issues will be the focus of our future work.
Ethic Considerations
In this article, we introduce a medical large language model, designated as MedCare, which incorporates several ethical considerations crucial for its deployment and usage in sensitive settings:
Performance vs. Potential Risks
While MedCare demonstrates enhancements over previous general and specialized medical models in knowledge reasoning and performance on downstream tasks, it’s important to acknowledge the inherent limitations of large language models. Notably, these models can exhibit "hallucinations" or generate misleading information. Additionally, the training datasets might harbor undiscovered biases, which could inadvertently influence the model’s outputs. Given these concerns, we emphasize that MedCare is not suitable for providing medical advice or for use in direct clinical applications.
Data Ethics and Privacy Compliance
The datasets employed in this study, including the knowledge testing data and the CBLUE dataset, are all publicly available and open-source and thus do not pose ethical dilemmas concerning their usage. However, the clinical data from CCTE used in training and testing, which involves report generation, image analysis, and discharge instructions, originates from hospital inpatient records. We have taken stringent measures to ensure the privacy and confidentiality of this information. All personal identifiers have been removed to maintain anonymity, ensuring no individual can be recognized from the data used. During the data collection, patients signed informed consent forms and were fully aware of the data usage methods described in this paper. Additionally, the usage of this data has been reviewed and approved by the corresponding hospital ethics committees. The specific approval numbers will be provided after the end of the review. This ensures that the data usage in this paper fully complies with ethical standards and privacy protection regulations.
Appendix A Related Works
Bilingual medical large language models
Large language models such as GPT-4 (OpenAI, 2022), PaLM (Chowdhery et al., 2023) and LLaMA (Touvron et al., 2023a, b) have achieved superior zero-shot performance across tasks and serve as interactive chatbots to interact with humans. However, trained on little medical-oriented data and Chinese data, these LLMs are not useful for medical conversations and consultations, especially for Chinese medical scenarios. Therefore, a lot of work has been done to fine-tune base models on medical data to obtain large medical models. Med-PaLM (Singhal et al., 2022) is grounded on Flan-PaLM (Chung et al., 2022; Chowdhery et al., 2023) to encode clinical knowledge. Med-PaLM2 (Singhal et al., 2023b) as a successor, reduce the gap with doctors by fine-tuning in the medical data and using modern prompting strategies. Apart from grounding on super-large language models such as PaLM or GPT, many works attempt to build medical-LLM on deployable sizes of LLMs, including 7B and 13B. DoctorGLM (Xiong et al., 2023), ChatGLM-Med (Wang et al., 2023b), and Bianque-2 (Chen et al., 2023b) are all built on ChatGLM to support acceptable bilingual medical consultation. Other work like MedicalGPT (Xu, 2023), Huatuo (Wang et al., 2023a) and HuatuoGPT-II (Chen et al., 2023a) are built on LLaMA-series and enlarge the vocabulary to support Chinese conversations.
Parameter efficient fine-tuning
Full fine-tuning effectively adapts base large language models to downstream tasks but also consumes significant computational resources with the increasing size of models and the number of tasks. To address this, Parameter-Efficient Fine-Tuning (PEFT) methods have been introduced. These methods freeze the base language models, modifying only a negligible number of parameters during the training phase, yet achieving similar or even superior performance with limited fine-tuning data. Among these methods, Adapter-Tuning (Rebuffi et al., 2017; Houlsby et al., 2019; Lin et al., 2020; Pfeiffer et al., 2021) was the pioneering architecture that connected two additional projection layers to the pretrained language model. In addition to incorporating additional modules, Prefix-tuning (Li and Liang, 2021) introduces learnable prefix tokens and prepends them before the input prompts. Differently, Prompt-Tuning (Lester et al., 2021) utilizes learnable prompt tokens for each task, as a multi-task PEFT approach in NLU scenarios. Following these, P-Tuning (Liu et al., 2023c) and P-Tuning-v2 (Liu et al., 2022b) move away from explicit prompts and employ a prompt-generator to convert pseudo prompts into task prompts, allowing decoder-only models to also perform NLU tasks. However, these methods introduce additional priors and significant inference latency. In contrast, Low-Rank Adaptation (LoRA)(Hu et al., 2022) and its variant, Weight-Decomposed Low-Rank Adaptation (DoRA)(Liu et al., 2024b), take a different approach. LoRA updates original parameters with two low-rank matrices without assuming any specific task or architecture, eliminating inference latency by merging back these two matrices to the original weight. DoRA extends this by incorporating weight decomposition, achieving performance comparable to full fine-tuning. Nonetheless, transferring these methods to multi-task learning scenarios without manual adjustments remains a challenge.
Type Task Description Size Metrics
MedQA Chinese Mainland Medical License Exams (USMLE) 3,425 Accuracy
MMedBnech Chinese subset of the Multilingual Medical Benchmark 3,425 Accuracy
CMB Comprehensive Multi-level Assessment for Medical Knowledge 11,200 Accuracy
CMExam Chinese National Medical Licensing Examination 6,811 Accuracy
CMMLU Chinese Massive Multitask Language Understanding 1,354 Accuracy
CEval A Multi-Level Multi-Discipline Chinese Evaluation 41 Accuracy
PLE Pharmacy Pharmacist Licensure Examination Pharmacy track 480 Accuracy
Knowledge-Intensive Tasks PLE TCM Pharmacist Licensure Examination Traditional Chinese Medicine track 480 Accuracy
Report Generation Analysis of Abnormal Indicators in Physical Examination Reports 50 GPT-4 Eval
Image Analysis Analysis of the Medical Image Reports 50 GPT-4 Eval
Discharge Instruction Providing patients with clear and comprehensive guidance 50 GPT-4 Eval
CCTE Examination Education Guide students in understanding the thought process behind examination questions 50 GPT-4 Eval
CMeEE Chinese Medical Named Entity Recognition 500 Micro-F1
CMeIE Chinese Medical Text Entity Relationship Extraction 600 Micro-F1
CHIP-CDN Clinical Terminology Normalization 600 Micro-F1
CHIP-CDEE Clinical Discovery Event Extraction 600 Micro-F1
IMCS-V2-NER Intelligent Medical Conversation System Named Entity Recognition 600 Micro-F1
CHIP-MDCFNPC Medical Dialog Clinical Findings Positive and Negative Classification 600 Micro-F1
IMCS-V2-SR Intelligent Medical Conversation System Symptom Recognition 600 Micro-F1
IMCS-V2-DAC Intelligent Medical Conversation System Dialogue Action Classification 800 Macro-F1
IMCS-V2-MRG Intelligent Medical Conversation System Medical Report Generation 600 RougeL
CHIP-CTC Clinical Trial Criterion 1100 Micro-F1
CHIP-STS Semantic Textual Similarity 600 Micro-F1
KUAKE-IR Information Retrieval 600 Micro-F1
KUAKE-QIC Query Intent Criterion 660 Micro-F1
KUAKE-QQR Query Query Relevance 600 Micro-F1
KUAKE-QTR Query Title Relevance 600 Micro-F1
CBLUE MedDG Medical Dialog Generation 600 RougeL
Table 5: Statistics of the evaluated medical tasks. $\dagger$" indicates that we only choose the questions related to the medical domain for the task.
Appendix B Experiments Details
B.1 Implementation Detail
For all sizes of MedCare, we set the batch size to 128 and fine-tuned the models with 1 epoch for the miscellaneous knowledge aggregation step. The learning rate is 2e-4, with a warmup_ratio=0.03 and the cosine learning schedule. The maximum length of the training sample is configured to 3072. We fine-tune the model with MKA stage for 1 epoch and DA stagê for 3 epochs. The orthogonal weight factor $\lambda=1$. For the configuration of the PEFT, the LoRA rank $r$ and $\alpha$ are fixed at 16 and 32, respectively. The number of the shared experts is set to 2, and the total number of mixture experts is set to 8, with 2 experts activated for each token during the training. We only adopt MoLoRA for the linear layers in the feed-forward network (FFN) blocks and adopt normal LoRA for the linear layers in the attention blocks. All the experiments are conducted on 8$\times$×A100 80G GPUs.
B.2 Baseline Models
We selected various models as baselines for comparing their performance on medical tasks. For the general open-sourced models, we choose Baichuan2-7B/13B-Chat (Yang et al., 2023), Qwen1.5-7B/14B-Chat (Bai et al., 2023), ChatGLM2/3-6b (Zeng et al., 2022), LLaMA2-7B/13B-Chat (Touvron et al., 2023a), and LLaMA3-8B-Instruct (AI, 2024). We choose HuatuoGPT-II (7B/13B) (Chen et al., 2023a) for the medical open-sourced models. Additionally, we also choose the ChatGPT (OpenAI, 2022) as the type of closed-source model with strong performance.
B.3 Fine-tuning Corpus
The corpus used for the MKA fine-tuning stage contains nearly 400k samples in total. For the publicly available parts of data, it contains 80k multiple-choice of the question-answering sample with rational from the training set of the MMedBench (Qiu et al., 2024) and CMExam (Liu et al., 2024a), and 68k samples from the PromptCBLUE (Zhu et al., 2023), which transfer CBLUE (Zhang et al., 2022a) into a pure text format using specific prompts, 30k multi-turn medical conversations sampled from the HuatuoGPT-sft-data-v1 (Wang et al., 2023a) and 50k single-turn question-answering data generated by GPT-4 from HuatuoGPT2_sft_instruct_GPT4_50K (Chen et al., 2023a). The private portion of the data contains about 200k training samples, with each of the four clinical tasks having 50k training samples each.
B.4 Testing Benchmarks
A detailed summary of the descriptions, quantities, and evaluation methods of all test data can be found in Table 5.
Medical Knowledge Exams
Medical licensing exams measure large language models’ medical knowledge and reasoning capabilities, serving as a common testing method. For this type of Benchmarks, we follow the experiments setting of Chen et al. (2023a). The examination benchmarks include the Chinese Mainland test set of MedQA (Jin et al., 2021), the Chinese test set of MMedBench (Qiu et al., 2024), and two comprehensive Chinese medical exam datasets, validation set of the CMB (Wang et al., 2023d) and the test set of the CMExam (Liu et al., 2024a). We also collect the medical parts of the general benchmarks, including the validation set of the CMMLU (Li et al., 2023a) and CEval (Huang et al., 2023). We also test the performance of the models on the flash exam questions from the 2023 Chinese National Pharmacist Licensure Examination, collected by Chen et al. (2023a).
Medical Alignment Tasks
The characteristic of medical alignment Tasks lies in their unique input and output format requirements, primarily evaluating the models’ ability to follow the instructions in the clinical scenario. We mainly choose two types of alignment tasks to evaluate the models. The first is CBLUE (Zhang et al., 2022a), a Chinese multi-task medical dataset encompassing 16 distinct medical NLP tasks. The second is the proposed Chinese Clinical Task Evaluation (CCTE), which comprises four clinical tasks: Report Generation, Image Analysis, Discharge Instruction, and Examination Education.
Appendix C Chinese Clinical Tasks Evaluation
C.1 Tasks Descriptions
Report Generation
In the Report Generation task, there are over 600 different types of test items and more than 3,600 test reports containing various combinations of these test items. The input comprises patient-specific laboratory test report data. The model is tasked with analyzing and identifying which of the patient’s test indicators deviate from normal ranges, providing insights into potential underlying causes and offering relevant recommendations for further action. Typically, the number of test items presented in the reports ranges from 2 to 16. To effectively accomplish this task, the model must demonstrate a robust capability for recognizing and interpreting extended contextual information, as well as possess an advanced understanding of medical knowledge. This is crucial for ensuring accurate analysis and the provision of actionable medical advice based on the test results. The data example is shown in Figure 7.
Figure 7: Case Example of Report Generation tasks.
Image Analysis
In the Image Analysis task, the primary focus is on generating, analyzing, and diagnosing based on descriptive reports derived from medical imaging. This task involves interpreting detailed reports associated with 14 types of medical image reports, including Rapid Pathology, Endoscopy, Magnetic Resonance Imaging, Electrocardiogram, Computed Tomography, Color Ultrasound, Digital Subtraction Angiography, Computed Radiography, Routine Pathology, Nuclear Medicine, Gastrointestinal Pathology, Endoscopy Examination, Radio Frequency, and Immunohistochemistry Pathology. The model must effectively parse and understand these textual descriptions to identify any noted abnormalities, correlate them with potential medical conditions, and provide diagnostic insights. The data example is shown in Figure 8.
Figure 8: Case Example of Image Analysis tasks.
Discharge Instruction
In the Discharge Instruction task, the primary objective is to generate comprehensive medical advice and instructions for patients at the time of their discharge based on their admission details, treatment processes, and discharge conditions. This task involves synthesizing information from a patient’s entire hospital stay, encompassing initial symptoms, diagnostic findings, treatments administered, and the patient’s response to those treatments. The model must adeptly process and integrate this wide array of medical data to formulate clear and precise discharge instructions. These instructions typically include guidelines on medication management, wound care, lifestyle adjustments, follow-up appointments, and signs of potential complications that should prompt immediate medical attention. The data example is shown in Figure 9.
Figure 9: Case Example of Discharge Instruction tasks.
Examination Education
The primary focus of the Examination Education task is on enhancing models’ ability to explain their decision-making processes, particularly in medical contexts. This task involves evaluating the model’s capability to provide detailed explanations and analyses of its responses. This not only aids in increasing the interpretability of large language models in healthcare applications but also serves as a valuable tool for medical students preparing for exams. For the data of the Examination Education, we sample the 50 samples from the test set of the MMedBench (Qiu et al., 2024) for each question in it contains the ground truth rational. The data example is shown in Figure 10.
Figure 10: Case Example of Examination Education tasks.
Figure 11: The prompt of the GPT-4 evaluation for CCTE tasks.
C.2 Evaluation Method
For the evaluation method of the CCTE dataset, we adopt the GPT4 evaluation to solve the problem that clinical tasks are difficult to evaluate. Following the evaluation pipeline of Wang et al. (2023d), we score each response across four aspects — Fluency, Relevance, Completeness, and Medical Proficiency — using a grading scale from 1 to 5. We input the original data from clinical data for this task, along with the model’s responses, into GPT-4 as evaluation criteria to prevent biases in scoring due to medical knowledge hallucinations in GPT-4. Simultaneously, GPT-4 leverages its capabilities to score language metrics such as fluency. The prompt of the GPT-4 is shown in Figure 11.
MODEL Pharmacist Licensure Examination (Pharmacy) Pharmacist Licensure Examination (TCM) AVERAGE
Optimal
Choice
Matched
Selection
Integrated
Analysis
Multiple
Choice
Total
Score
Optimal
Choice
Matched
Selection
Integrated
Analysis
Multiple
Choice
Total
Score
Llama2-7B 16.88 24.09 15.00 0.00 18.54 15.63 17.27 21.67 2.50 16.04 17.29
Llama2-13B 15.63 26.82 15.00 0.00 19.38 18.13 23.18 21.67 0.00 19.38 19.38
Llama3-8B 41.25 42.72 30.00 15.00 38.33 38.13 30.90 35.00 27.50 33.54 35.94
ChatGLM2-6B 37.00 36.80 25.00 31.70 35.30 33.10 37.30 35.00 37.30 33.70 34.50
ChatGLM3-6B 39.50 39.10 10.50 0.20 34.60 31.80 38.20 25.00 20.00 32.90 33.75
Biachuan2-7B 51.20 50.90 30.00 2.60 44.60 48.10 46.00 35.00 7.50 42.10 43.35
Biachuan2-13B 43.80 52.70 36.70 7.90 44.20 41.30 46.40 43.30 15.00 41.70 42.95
Qwen1.5-7B 29.38 37.73 30.00 17.50 32.29 30.63 34.09 18.33 12.50 29.17 30.73
Qwen1.5-14B 53.75 51.36 46.67 12.50 48.33 40.00 37.73 43.33 27.50 38.33 43.33
ChatGPT 45.60 44.10 36.70 13.20 41.20 34.40 32.30 30.00 15.00 31.20 36.20
HuatuoGPT-II-7B 41.90 61.00 35.00 15.70 47.70 52.50 51.40 41.70 15.00 47.50 47.60
HuatuoGPT-II-13B 47.50 64.10 45.00 23.70 52.90 48.80 61.80 45.00 17.50 51.60 52.25
GPT-4 66.88 65.91 46.67 40.00 61.67 39.38 50.45 45.00 32.50 44.58 53.13
MedCare-1.8B 33.75 42.73 18.33 0.25 33.33 40.00 40.91 31.67 12.50 37.08 35.21
      $\mapsto$ w/o DA 34.38 44.55 30.00 5.00 36.04 40.00 42.73 45.00 12.50 39.58 37.81
MedCare-7B 51.88 53.18 40.00 17.50 48.13 48.75 59.09 51.67 25.00 51.88 50.00
      $\mapsto$ w/o DA 50.63 60.91 38.33 15.00 50.83 52.50 62.27 38.33 30.00 53.33 52.08
MedCare-14B 53.75 60.91 40.00 20.00 52.50 51.88 58.64 48.33 32.50 52.92 52.71
      $\mapsto$ w/o DA 59.38 67.27 36.67 25.00 57.29 56.88 57.73 51.67 27.50 54.17 55.73
Table 6: Results of the 2023 Chinese National Pharmacist Licensure Examination. It consists of two separate Examinations including Pharmacy track and Traditional Chinese Medicine (TCM) Pharmacy track. The results of the baseline models are obtained from Chen et al. (2023a)